Ticket Number: HEA-NOC/20080207-1 Ticket Status: CLOSE
Ticket Type: unscheduled Resolver: HEA-NOC
Ticket Opened: 20080207 09:23 UTC Problem Start: 20080206 22:25 UTC
Ticket Closed: 20080208 20:01 UTC Problem End: 20080207 15:58 UTC
Site/line: St. Angela's
The connection was lost to St Angela's College at 22:25 on the 6th Feb.
No traffic to or from the site.
20080207 09:19 BN No contact could be made with a technical contact on
site, this will be retried soon. Reported fault to BT Ireland, they are
investigating with reference number 1-557125317.
20080207 11:17 BN BT Ireland investigations are ongoing. They have
informed us that they are going to swap out a card that forms part of the
circuit, but they have no ETA for this work.
20080207 16:07 BN Further BT testing indicated that the line is clear.
Made contact with St Angela's and reloaded the on-site HEAnet equipment.
After a reload the link was restored. Will continue to monitor.
20080208 20:01 BN Link has remained stable since reload.
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20080207-1
HEAnet Limited. Registered in Ireland, No. 275301.